A 36.8 g/l solution of K4[Fe(CN)6] is isotonic with 5% solution of non volatile solute find molecular mass of solute 
 

Grade:12

1 Answers

Rituraj Tiwari
3 years ago
Degree of dissociation, x=50%=0.5

Van't Hoff factor, i=

total number of moles initially

number of moles after dissociation

=

1

1+4x

i=1+4(0.5)=1+2=3

Decimolar solution means the molarity of solution is M/10 or 0.1 M. That is, 0.1 mole of solute is dissolved in 1 L of solution. Since, 1 L=1/1000 m

3

, therefore it can be said, 0.1 mole of solute is dissolved in 1/1000 m

3

volume of solution.

We have, πV=i nRT

π=

V

n

RT

π=osmotic pressure

V=volume of solution

n=number of moles of solute

R=universal gas constant=8.314 Nm mol

−1

K

−1

T=temperature=300K

i=Van



t Hoff factor=3

π=3×

0.001

0.1

×8.314×300

π=7.48×10

5

Nm

−2

=

1.013×10

5

7.48×10

5

atm

π=7.38 atm